OpenStudy (anonymous):

Matthew opens a store credit card when he purchases a new sound system. The interest rate is 22.45%. He charges $4,120 to the card, and can pay $1,450 per month. What will the total cost of his purchase be?

OpenStudy (espex):

First off, that interest rate is outrageous! Second, we are assuming a simple interest formula so the approach you want to take is to add 22.45% of each month's total and subtract your monthly payment. When you get to 0 you will have the cost of each month plus the amount of interest paid each month, for your total purchase price.

OpenStudy (anonymous):

I got 5,044.94 is that right?

OpenStudy (anonymous):

?

OpenStudy (espex):

I got 5302 but am going back over to see if I made an error.

OpenStudy (espex):

Again I got 5302.51, but I assumed he made his payments on time so only the residual would be charged interest. 4120-1450 = 2670 2670 + 599.415 = 3269.415 3269.415 - 1450 = 1819.415 1819.415 + 408.459 = 2227.877 2227.877-1450= 777.874 777.874 + 174.633 = 952.506 So three payments of 1450 + a residual of 952.506 = 5302.506 rounded to 5302.51

OpenStudy (espex):

Something about this just seemed wrong, so I did some searching and concluded that this is the appropriate answer, maybe @myininaya or @amistre64 or @hero could check it. To start, 22.45% was probably an annual rate, meaning that it was 0.018708333/mo To find the number of payments necessary I found this formula: N = −log(1−iA/P) / log(1+i) Where N is number of payments, i is the interest rate per month, A is the payment amount P is the amount borrowed. \[\large \frac{-\log_{10}(1-\frac{4120*0.018708333}{1450})}{\log_{10}(1+0.01870833) }\] equating to 2.9469 payments. \[2.9469 * 1450 = 4273.03\]

OpenStudy (anonymous):

Flint purchased a boat for $10,815. He made a down payment of $1,175. He applied for a six-year installment loan with an interest rate of 9.4% in the amount of $9,640. What is the total cost of the boat after six years? could you help me with this one its given me much trouble

OpenStudy (amistre64):

Matthew opens a store credit card when he purchases a new sound system. The interest rate is 22.45%. He charges $4,120 to the card, and can pay $1,450 per month. What will the total cost of his purchase be? \[B_n = B_{n-1}(.2245)-1450\] when Bn = 0 or less, we would have the nth payment and can calculate the overall monies paid is what im thinking

OpenStudy (amistre64):

B0 = 4120 B1 = 4120(.2245) - 1450 B2 = (4120(.2245) - 1450)(.2245) - 1450 = 4120(.2245)^2 - 1450(.2245) - 1450 B3 = 4120(.2245)^3 - 1450(.2245)^2 - 1450(.2245) - 1450 = 4120(.2245)^3 - 1450 (.2245^2 + .2245 + 1) B4 = 4120(.2245)^4 - 1450 (.2245^3 + .2245^2 + .2245 + 1) \[B_n=4120(.2245)^n -1450(.2245^{n-1} + .2245^{n-2}+... + .2245 + 1)\] \[B_n=4120(.2245)^n -1450\frac{1-.2245^{n}}{1-.2245}\] right?

OpenStudy (espex):

You're taking a summation approach..

OpenStudy (amistre64):

yes, a recursive approach is how i might surmise it :)

OpenStudy (amistre64):

when Bn = 0 or less we would have the nth payment and can deduce the amount from that in my idea

OpenStudy (espex):

For @weaver125 's second question I was thinking i=prt, where i is the interest paid, p is the principle amount, r is the rate and t is the term of the loan. Add the interest to the loan amount plus the down.

OpenStudy (espex):

Initially when I looked at this problem I imagined it as one that a simple i=prt or some other simplistic interest formula would handle. I honestly did not expect it to go this far. :)

OpenStudy (amistre64):

i = prt is simple interest

OpenStudy (espex):

Right, you are offered no payments (for the second question) just a simple "total". Loan amount + interest paid + down payment = total is what I was thinking.

OpenStudy (amistre64):

without payments, we might be able to assume that total compounded at the end of the time limit is the amount to be paid ...

OpenStudy (amistre64):

Flint purchased a boat for $10,815. He made a down payment of $1,175. He applied for a six-year installment loan with an interest rate of 9.4% in the amount of $9,640. What is the total cost of the boat after six years? 1175 + 9640(1.094)^6 is my guess

OpenStudy (espex):

That is exactly what I was thinking. :)